Francesca traveled for 5h in her car at a constant speed of 90km/h and then traveled in a bus at a constant speed for 3 h. If the total distance she traveled was 795 km, what was the speed of the bus?
Answer by checkley71(8403)

5*90+3S=795
450+3S=795
3S=795-450
3S=345
S=345/3
S=115 KM IS THE SPEED OF THE BUS.